40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
    Reply  Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating 
       Reply  Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed SRMoplevKWMtune.png
    Reply  Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit ITMXoplev.pngITMXoplevservo_ON.png
       Reply  Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd ITMXoplev1103p#2.pngITMXoplev.png
          Reply  Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm ETMX-oplev.jpg
             Reply  Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum ETMXoplev2lens.png
             Reply  Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev ETMXoplevlensAR.png
Message ID: 5694     Entry time: Wed Oct 19 10:49:35 2011     In reply to: 5636
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: SRM oscillation removed 

Quote:

The SRM oplevs were found to be oscillating because of a small phase margin.

I reduced the gains to the half of them. The peak which Steve found today maybe due to this oscillation.

Quote from #5630

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it. 

 Kiwamu removed the 18.3 Hz ocsillation by turning down the oplev servo gain.

Attachment 1: SRMoplevKWMtune.png  52 kB  | Hide | Hide all
SRMoplevKWMtune.png
ELOG V3.1.3-